OpenStudy (anonymous):

Factor completely: n^2-9n+8 Please help.

OpenStudy (unklerhaukus):

what are some factors of 8, that add to -9?

OpenStudy (unklerhaukus):

(the factors can be positive or negative)

OpenStudy (anonymous):

-8 and -1 ?

OpenStudy (unklerhaukus):

good

OpenStudy (anonymous):

(n-8)(n-1) n=8 or n=1 am I right ? @UnkleRhaukus

OpenStudy (unklerhaukus):

your first line is all this question is asking for if it was n^2-9n+8=0 then your would be right that x=8 or 1 but there was no equal sign in the original expression

OpenStudy (anonymous):

I know, but the whole point of this question is I think to find what n could be..

OpenStudy (unklerhaukus):

if you want to check it just expand the brackets, you should get the expression in the question (n-8)(n-1)=(n×n)+(n×-1)+(-8×n)+(-8×-1)

OpenStudy (unklerhaukus):

oh i thought you just had to Factor completely

OpenStudy (anonymous):

oh... I see. thank you;D

OpenStudy (unklerhaukus):

but if this was just a sub question to a larger problem , then your probably right

OpenStudy (anonymous):

ohh.... could you help my with this one too: 7x^2-45-28 what should I do with the 7? @UnkleRhaukus

OpenStudy (unklerhaukus):

this one isn't so simple i would set it equal to zero , and apply the quadratic formula \[ax^2+bx+c=0\] \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[ax^2+bx+c=(x-x_1)(x-x_2)\]

OpenStudy (anonymous):

but they want me to factor it Completely...

OpenStudy (unklerhaukus):

(x−x_1)(x−x_2) is the factored form

OpenStudy (unklerhaukus):

7x^2-45-28=0 what is a,b,c?